EV007A

PI: Venturi
Title: 1334-127,1504-166,2243-123

Session: June 1996 (session 2)
Frequency: 6 cm
Mode: E
Obs time (UT): 167 EU 14:00(15/06)-07:00(16/06)
Stations: JbNtOnWbToSmUrSh
Processor: Bonn
EU Mk3 Tapes/St.: 3

Comment


To give feedback on this experiment, click here

Experiment Feedback

Onsala (John Conway at Sun Jun 16 03:32:02 EDT 1996) Success
No known problems

Torun (Roman Feiler at Sun Jun 16 08:07:05 EDT 1996) Minor failures
several scans not recorded properly due to tracking problems

Noto (Corrado Trigilio at Tue Jun 18 09:21:43 EDT 1996) Minor failures
The first two scans of the second tape have been partially lost for problems during tape change (10 minuts in total). No other problems reported

Westerbork (Tony Foley at Thu Jun 20 05:49:30 EDT 1996) Minor failures
Some scans were at too low an elevation according to our online system (recorded anyway) at 16:28 to 16:55 UT.

Jodrell (David Henstock at Fri Jun 21 08:58:58 EDT 1996) Minor failures
Continual "tape head 2 not in requested position errors" during experiment, but as this is read head it should not affect data. Some of scan 19:19-19:30 UT lost due to problem with MkII telescope tracking in elevation. Field system crashed @ 6:30 UT and therefore some or all of the last half hour of data will have been lost

Urumqi (Lee Suqin at Tue Jul 2 01:21:04 EDT 1996) Failure
On day 167 It was rain. The comunication between the tape-recorder and the computer was not good. The data maybe incorrect. Failed completely.

Simeiz (Alexandre Volvatch at Tue Jul 2 06:09:56 EDT 1996) Minor failures
4 passes (22:21-23:16) were omited because El of source was very low (El<5deg).

Shanghai (Xiao Yu Tan at Mon Jul 15 12:45:56 EDT 1996) Minor failures
Clock offset was -0.6 micro sec. All obs took place with some failures and only LCP is avilable. Following data is invalid because of tape recoeder vacuum trouble: 14:06:44 to 14:13:04; 14:16:00 to 14:22:20; 14:36:43 to 14:43:04; 14:52:43 to 14:59:04; 15:06:43 to 15:13:04; 15:20:40 to 15:27:04. Weather was overcast.

HartRAO (Jonathan Quick at Wed Jul 17 08:05:52 EDT 1996) Success
Phase Cal level may have been too high during this expt. Tapes only arrived on 25 July 1996 No other problems to report.

Calibration (Fredrik T Rantakyro at Thu Oct 3 11:53:47 EDT 1996) Minor failures
An ANTAB file for this experiment can be found on VLBEER in the JUN96 directory.